OpenStudy (anonymous):

If sin x = .21,what is cos x?

OpenStudy (amistre64):

one way to solve this would be to use the pythagorean identity of sin and cos: sin^2 + cos^2 = 1

OpenStudy (anonymous):

sin^2x+c0s^2x=1 sin^2x=1-cos^2x (.21)^2=1-cos^2x 0.0441=1-cos^2x cos^2x=1-0.o441 cos^2x=0.9559 cosx=+-0.9777